OpenStudy (anonymous):

A segment with endpoints I (5, 2) and J (9, 10) is divided by a point K such that IK and IJ form a 2:3 ratio. Find the y value for K. 4.6 5.4 4.8 5.2

OpenStudy (anonymous):

i need some help pronto

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@gswag98

OpenStudy (anonymous):

im not good at this

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@AkashdeepDeb do you know how to do this?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

how did you get that melani?

OpenStudy (anonymous):

i dont understand at all

OpenStudy (akashdeepdeb):

Baahh! i was going to type an answer and I refreshed the page. This is known as internal division. The point k divides the line into 5 parts. 2 parts with one side and 3 parts with the other. Check this link: It explains it well: http://www.teacherschoice.com.au/Maths_Library/Analytical%20Geometry/AnalGeom_3.htm

OpenStudy (anonymous):

@AkashdeepDeb could you use the formula and show me how i would find it idk how to use that

OpenStudy (anonymous):

that formula is really hard for me

OpenStudy (anonymous):

im not asking for an answer just if you could plug in the numbers in the equation for me

OpenStudy (phi):

find the difference (in the y values) from point I to point J divide that difference by 5

OpenStudy (phi):

multiply by 2 add that to the starting y value (the y value of point I)

OpenStudy (akashdeepdeb):

No problem. It is very simple. Here is a general example.  \[(x,y) = (\frac{mx_2 + nx_1}{m+n},\frac{my_2 + ny_1}{m+n})\] \[x = \frac{mx_2+nx_1}{m+n}\]\[y = \frac{my_2 + ny_1}{m+n}\] Basically, the x value of the point dividing the line uses the x values of the 2 end points. And the y value of the point dividing the line uses the y values of the 2 end points. Getting this? :)

OpenStudy (anonymous):

this is hard im only in tenth man and i failed geometry

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

is there a simpler way idk how to do this man its really confusing

OpenStudy (anonymous):

ive never been taught on how to use that

OpenStudy (akashdeepdeb):

I understand. Don't be overwhelmed by what i wrote up there. You most probably would understand it in a few months, when you may do it in school. I am not really aware of any other method. [Other than graphing it or using the formula I wrote above]. But there is always another way. Try tagging someone else, who you think could help. :)